OpenStudy (anonymous):

The length of a room is twice the breadth of the room. The cost of carpeting it at Rs. 1 per square metre is Rs. 1800 and the cost of white washing it at 50 paise per square metre is Rs. 265, excluding the area of doors and windows of area 10 square metres. The height of the room is

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

bcoz The cost of carpeting it at Rs. 1 per square metre is Rs. 1800 hence the area of the floor must be 1800 and since lenght is twice breadth so we can say that square of breadth is 1800/2 can you proceed now (altrenately use breadth x and length 2x and 2x^2 =1800)

OpenStudy (tkhunny):

Please don't tag the world. We're either here or we're not. Multiple tags are more annoying than they are helpful. You haven't shown any work. What have you done? How far have you gotten? Define variables for missing information. L = Length of floor B = Breadth of floor H = Height of Room Do your best to translate the language to formulas. "The length of a room is twice the breadth of the room" L = 2*B "The cost of carpeting it at Rs. 1 per square metre is Rs. 1800" L*B = 1800 "the cost of white washing it at 50 paise per square metre is Rs. 265, excluding the area of doors and windows of area 10 square metres." 2*L*H + 2*B*H - 10 <== There's the area of the walls. Now what?

OpenStudy (anonymous):

sorry network problem.i found length and breadth.h=1/180

OpenStudy (anonymous):

h=1/18

OpenStudy (aum):

I am continuing from tkhunny's reply: So far we have: L = 2B ---- (1) LB = 1800 ----- (2) Two equations and two unknowns so far and so we can solve for L and B. Put L = 2B from (1) into (2): 2B * B = 1800 B^2 = 900 B = 30 m and from (1): L = 60 m "cost of white washing it at 50 paise (i.e., Rs. 0.50) per square metre is Rs. 265" Area white washed = 265 / 0.5 = 265*2 = 530 m^2. 4 walls need to be white washed and I am assuming the ceiling too. If H is the height of the room, then the area of the 4 walls + ceiling = LH + BH + LH + BH + LB = 2LH + 2BH + LB "excluding the area of doors and windows of area 10 square metres." 2LH + 2BH + LB - 10 = 530 m^2 ---- (3) Put L = 60, B = 30 in (3) and solve for H

OpenStudy (aum):

Forget the ceiling in the above equation as it does not yield proper value for H. Rewrite equation (3) as: 2LH + 2BH - 10 = 530 m^2 ---- (3) or LH + BH - 5 = 265 H(L + B) = 270 H(60+30) = 270 H * 90 = 270 H = 270 / 90 = 3 m. The height of the room is 3 meters (or 3 metres as spelled in your country).

OpenStudy (anonymous):

i have another question.i want to read geomtry and trigonometry.i many question unsolved

OpenStudy (aum):

Close this question and post it as a separate question, one question per posting.